### v5

# Characterizing categoricity in the class $\mathrm{Add}(M)$

Abstract
We show that the condition of being categorical in a tail of cardinals can be characterized for the class of $R$-modules of the form $\mathrm{Add}(M)$. More precisely, let $R$ be a ring and $M$ be an $R$-module which can be generated by $\leq\aleph$ elements. Then $\mathrm{Add}(M)$ is $\kappa$-categorical in all $\kappa>\|R\|+\aleph+\aleph_{0}$ if and only if $\mathrm{Add}(M)$ is $\kappa$-categorical in some $\kappa>\|R\|+\aleph+\aleph_{0}$; if and only if every $R$-module of cardinal $\kappa$ in $\mathrm{Add}(M)$ is $M$-free for all $\kappa>\|R\|+\aleph+\aleph_{0}$; if and only if every $R$-module of cardinal $(\|R\|+\aleph+\aleph_{0})^{+}$ in $\mathrm{Add}(M)$ is $M$-free.
As an application, we show that the class of pure-projective $R$-modules is categorical in some (all) big cardinal if and only if $R$ is $R\cong M_{n}(D)$ where $D$ is a division ring and $n\geq 1$; the class of semisimple $R$-modules is categorical in some (all) big cardinal if and only if $R$ admits a unique simple module up to isomorphism, partly answering a question proposed in [5, Mazari-Armida M., Characterizing categoricity in several classes of modules. J. Algebra 617, 382-401 (2023)].
Key Words: $\mathrm{Add}(M)$; pure-projecitve modules, categoricity, Shelah’s Categoricity Conjecture.
2020 Mathematics Subject Classification: 16B70.
1. Introduction
[…]
The notion of $\lambda$-categoricity of some structure, which requires that there is a structure of size $\lambda$ and any two structures of size $\lambda$ are isomorphic, was started by Łoś [4] in 1954. It has played a key role in the development of model theory. Morley’s Categoricity Theorem [6], which asserts that if the class of structures of a first-order theory in a countable language is $\lambda$-categorical in some uncountable cardinal $\lambda$ then it is $\mu$-categorical in all uncountable cardinals $\mu$, can be seen as the birth of modern model theory. Currently, Shelah’s Categoricity Conjecture [8], which states that if $\mathbf{K}$ is an abstract elementary class which is $\lambda$-categorical in some cardinal $\lambda\geq\beth_{(2^{{\rm LS}(\mathbf{K})})^{+}}$, then $\mathbf{K}$ is $\mu$-categorical in all cardinals $\mu\geq\beth_{(2^{{\rm LS}(\mathbf{K})})^{+}}$, is a far-reaching generalization to Morley’s Categoricity Theorem.
Shelah’s Categoricity Conjecture is a long-standing open problem of abstract model theory for the general abstract elementary classes. It will be interesting to verify it for a concrete algebraic structure, such as a class of modules, being categorical in a tail of cardinals. Recently, Marcos Mazari-Armida [5] studied several classical classes of modules, such as the whole modules, absolutely pure modules, flat modules and locally (pure-)injective modules, in terms of categoricity. subsequently, Trlifaj [9] firstly characterized the categoricity of projective modules, and then showed that if $\mathcal{D}$ is a deconstructible class of modules, then the categoricity of $\mathcal{D}$ in a big cardinal is equivalent to its categoricity in a tail of cardinals. More accurate expression in cardinals can also be found in [7, Theorem 4.1].
In this paper, we will investigate the categoricity of the classes of the form $\mathrm{Add}(M)$, all direct summands of direct sums of copies of a given $R$-module $M$. More precisely, we obtain the following result.
Theorem 1.1.
(((=Theorem 2.2))) Let $R$ be a ring and $M$ be an $R$-module which can be generated by $\leq\aleph$ elements. Then the following statements are equivalent:
  1. (1)
    $\mathrm{Add}(M)$ is $\kappa$-categorical in all $\kappa>\|R\|+\aleph+\aleph_{0}$.
  2. (2)
    $\mathrm{Add}(M)$ is $\kappa$-categorical in some $\kappa>\|R\|+\aleph+\aleph_{0}$.
  3. (3)
    every $R$-module of cardinal $\kappa$ in $\mathrm{Add}(M)$ is $M$-free for all $\kappa>\|R\|+\aleph+\aleph_{0}$.
  4. (4)
    every $R$-module of cardinal $(\|R\|+\aleph+\aleph_{0})^{+}$ in $\mathrm{Add}(M)$ is $M$-free.
As applications, we can characterize the categorical property of semisimple modules and pure-projective $R$-modules in terms of ring theory.
Theorem 1.2.
(((=Theorem 2.5))) Let $R$ be a ring, $S\SS$S the class of all semisimple $R$-modules and $S$ the direct sum of all simple modules non-isomorphic to each other. Then the following statements are equivalent:
  1. (1)
    $S\SS$S is $\kappa$-categorical in all $\kappa>\|R\|+\aleph_{0}$.
  2. (2)
    $S\SS$S is $\kappa$-categorical in some $\kappa>\|R\|+\aleph_{0}$.
  3. (3)
    Every semisimple module of cardinal $\kappa$ is $S$-free for all $\kappa>\|R\|+\aleph_{0}$.
  4. (4)
    Every semisimple module of cardinal $(\|R\|+\aleph_{0})^{+}$ is $S$-free.
  5. (5)
    $R$ admits a unique simple module up to isomorphism.
Theorem 1.3.
(((=Theorem 2.7))) Let $R$ be a ring and $P\mathcal{P}roj$ the class of all pure-projective $R$-modules. Then the following statements are equivalent:
  1. (1)
    $P\mathcal{P}roj$ is $\kappa$-categorical in all $\kappa>\|R\|+\aleph_{0}$.
  2. (2)
    $P\mathcal{P}roj$ is $\kappa$-categorical in some $\kappa>\|R\|+\aleph_{0}$.
  3. (3)
    every pure-projective module of cardinal $\kappa$ is free for all $\kappa>\|R\|+\aleph_{0}$.
  4. (4)
    every pure-projective module of cardinal $(\|R\|+\aleph_{0})^{+}$ is free.
  5. (5)
    $R\cong M_{n}(D)$ where $D$ is a division ring and $n\geq 1$.
2. main results
We begin with the following famous result, which is due to Walker:
Lemma 2.1.
[10, Theorem 4.2] Let $\kappa$ be an infinite cardinal. If an $R$-module $M$ is a direct sum of $(\leq\kappa)$-generated submodules, then so is every direct summand of $M$.
We say a class of $R$-modules is $\lambda$-categorical in some cardinal $\lambda$ if there is an $R$-module of size $\lambda$ in the class and if any two $R$-modules of size $\lambda$ in the class are isomorphic. A class of modules is categorical in a tail of cardinals if there is a cardinal $\mu$ such that the class is $\lambda$-categorical in all $\lambda>\mu.$
Let $M$ be an $R$-module. Set $\mathrm{Add}(M)$ to be the class of all $R$-modules which is a direct summand of direct sums of copies of $M$. We call an $R$-module $N$ is $M$-free if $N\cong M^{(\kappa)}$ for some cardinal $\kappa.$ The following main theorem shows that the categoricity of $\mathrm{Add}{M}$ in a big cardinal is equivalent to its categoricity in a tail of cardinals.
Theorem 2.2.
Let $R$ be a ring and $M$ be an $R$-module which can be generated by $\leq\aleph$ elements. Then the following statements are equivalent:
  1. (1)
    $\mathrm{Add}(M)$ is $\kappa$-categorical in all $\kappa>\|R\|+\aleph+\aleph_{0}$.
  2. (2)
    $\mathrm{Add}(M)$ is $\kappa$-categorical in some $\kappa>\|R\|+\aleph+\aleph_{0}$.
  3. (3)
    Every $R$-module of cardinal $\kappa$ in $\mathrm{Add}(M)$ is $M$-free for all $\kappa>\|R\|+\aleph+\aleph_{0}$.
  4. (4)
    Every $R$-module of cardinal $(\|R\|+\aleph+\aleph_{0})^{+}$ in $\mathrm{Add}(M)$ is $M$-free.
Proof.
(3)(1)(2)(3$\Rightarrow$(1$\Rightarrow$(2)( 3 ) ⇒ ( 1 ) ⇒ ( 2 ) and (3)(4)(2)(3$\Rightarrow$(4$\Rightarrow$(2)( 3 ) ⇒ ( 4 ) ⇒ ( 2 ): Trivial.
(2)(3)(2$\Rightarrow$(3)( 2 ) ⇒ ( 3 ) Suppose $\mathrm{Add}(M)$ is $\kappa_{0}$-categorical in some fixed cardinal $\kappa_{0}>\|R\|+\aleph+\aleph_{0}$. We will show every $R$-module of cardinal $\kappa$ in $\mathrm{Add}(M)$ is $M$-free by dividing into two cases:
Case I: $\kappa\geq\kappa_{0}.$ Let $P$ be an $R$-module with cardinal $\kappa$ in $\mathrm{Add}(M)$. Then $P=\bigoplus\limits_{i\in I}P_{i}$ where each $P_{i}$ is $(\leq\kappa)$-generated $R$-module by Lemma 2.1. Since $\|P\|=\kappa>\|R\|+\aleph+\aleph_{0}$ and $\|P_{i}\|\leq\|R\|+\aleph+\aleph_{0}$ for each $i\in I$, we have $|I|=\kappa.$ Then we can give a partition $\{I_{\lambda}\mid\lambda\in\Lambda\}$ of $I$ with each $|I_{\lambda}|=\kappa_{0}$. Set $\bigoplus\limits_{i\in I_{\lambda}}P_{i}=P^{\prime}_{\lambda}$. Then $P=\bigoplus\limits_{\lambda\in\Lambda}P^{\prime}_{\lambda}$. Note that $\|P^{\prime}_{\lambda}\|=\kappa_{0}$. So $P^{\prime}_{\lambda}\cong M^{(\kappa_{0})}$ by (2)(2)( 2 ) as $\|M^{(\kappa_{0})}\|=\kappa_{0}$. Consequently, $P\cong\bigoplus\limits_{\lambda\in\Lambda}M^{(\kappa_{0})}\cong M^{(\kappa)}$.
Case II: $\|R\|+\aleph+\aleph_{0}<\kappa<\kappa_{0}.$ Let $P$ be an $R$-module with cardinal $\kappa$ in $\mathrm{Add}(M)$. Again, $P=\bigoplus\limits_{i\in I}P_{i}$ where each $P_{i}$ is $(\leq\kappa)$-generated $R$-module by Lemma 2.1. Since $\|P\|=\kappa>\|R\|+\aleph+\aleph_{0}$ and $\|P_{i}\|\leq\|R\|+\aleph+\aleph_{0}$ for each $i\in I$, we have $|I|=\kappa.$ Note that $\|M^{(\kappa)}\|=\kappa$. So $P\cong M^{(\kappa)}$ by (2).(2).( 2 ) . That is,
$P=\bigoplus\limits_{i\in I}P_{i}=\bigoplus\limits_{i\in J}Q_{i}$
with each $Q_{i}$ isomorphic to $M$, where $|J|=\kappa$. It is easy to see that $P$ is $M$-free by the following claim, as we will have $P=\bigoplus\limits_{\lambda\in\Lambda}(\bigoplus\limits_{j\in J_{\lambda}}Q_{j})\cong M^{(\kappa)}$.
Claim: There is a partition $\{I_{\lambda}\mid\lambda\in\Lambda\}$ (((resp., $\{J_{\lambda}\mid\lambda\in\Lambda\})$ of $I$ (((resp., $J)$ with $|I_{\lambda}|\leq\kappa$ (((resp., $|J_{\lambda}|\leq\kappa)$ for each $\lambda\in\Lambda$ such that
$\bigoplus\limits_{i\in I_{\lambda}}P_{i}=\bigoplus\limits_{j\in J_{\lambda}}Q_{j}$
for each $\lambda\in\Lambda.$
The rest of this proof is devoted to prove the above Claim. Suppose $i_{0}\in I$. We will first show that there exist subsets $I^{\prime}\subseteq I$ and $J^{\prime}\subseteq J$ of cardinal $\leq\kappa$ such that $i_{0}\in I^{\prime}$ and $\bigoplus\limits_{i\in I^{\prime}}P_{i}=\bigoplus\limits_{j\in J^{\prime}}Q_{j}$. Indeed, we first construct sets $I^{\prime}_{n}$ and $J^{\prime}_{n}$ of cardinal $\leq\kappa$ for each $n\in\mathbb{N}$ by induction as follows:
Start step: Set $I^{\prime}_{0}=\{i_{0}\}$ and $J^{\prime}_{0}=\emptyset$.
Inductive step: Suppose $I^{\prime}_{n}$ and $J^{\prime}_{n}$ have been defined. Since $(\bigoplus\limits_{i\in I^{\prime}_{n}}P_{i})+(\bigoplus\limits_{j\in J^{\prime}_{n}}Q_{j})$ is $(\leq\kappa)$-generated, so there exists a subset $J^{\prime}_{n+1}\subseteq J$ of cardinal at most $\kappa$ such that $(\bigoplus\limits_{i\in I^{\prime}_{n}}P_{i})+(\bigoplus\limits_{j\in J^{\prime}_{n}}Q_{j})\subseteq\bigoplus\limits_{j\in J^{\prime}_{n+1}}Q_{j}$. Similarly, since $(\bigoplus\limits_{i\in I^{\prime}_{n}}P_{i})+(\bigoplus\limits_{j\in J^{\prime}_{n+1}}Q_{j})$ is $(\leq\kappa)$-generated, there exists a subset $I^{\prime}_{n+1}\subseteq I$ of cardinal at most $\kappa$ such that $(\bigoplus\limits_{i\in I^{\prime}_{n}}P_{i})+(\bigoplus\limits_{j\in J^{\prime}_{n+1}}Q_{j})\subseteq\bigoplus\limits_{i\in I^{\prime}_{n+1}}P_{i}$.
Then, it is obvious that $I^{\prime}=\bigcup_{n\geq 0}I^{\prime}_{n}$ and $J^{\prime}=\bigcup_{n\geq 0}J^{\prime}_{n}$ have the required property.
Next, we define a chain of subsets $K_{0}\subseteq K_{1}\subseteq\cdots\subseteq K_{\lambda}\subseteq\cdots$ of $I$ and a chain of subsets $L_{0}\subseteq L_{1}\subseteq\cdots\subseteq L_{\lambda}\subseteq\cdots$ of $J$ for each ordinal $\lambda$ by transfinite induction as follows:
Start step: Set $K_{0}=L_{0}=\emptyset$.
limit ordinal step: If $\lambda$ is a limit ordinal set $K_{\lambda}=\bigcup_{\mu<\lambda}K_{\mu}$ and $L_{\lambda}=\bigcup_{\mu<\lambda}L_{\mu}$.
success ordinal step: For every ordinal $\mu$ such that $K_{\mu}=I$, set $K_{\mu+1}=K_{\mu}$ and $L_{\mu+1}=L_{\mu}$. Otherwise, if $K_{\mu}\subsetneq I$, choose an element $i_{0}\in I\setminus K_{\mu}$. By the claim there exist $I^{\prime}\subseteq I$ and $J^{\prime}\subseteq J$, both of cardinal $\leq\kappa$, such that $i_{0}\in I^{\prime}$ and $\bigoplus\limits_{i\in I^{\prime}}P_{i}=\bigoplus\limits_{j\in J^{\prime}}Q_{j}$. In this case set $K_{\mu+1}=K_{\mu}\cup I^{\prime}$ and $L_{\mu+1}=L_{\mu}\cup J^{\prime}$.
Finally, we are ready to prove the Claim. We obviously have $\bigoplus\limits_{i\in K_{\lambda}}P_{i}=\bigoplus\limits_{j\in L_{\lambda}}Q_{j}$ for every $\lambda$, and there exists an ordinal $\lambda^{\prime}$ such that $K_{\lambda^{\prime}}=I$. Then $L_{\lambda^{\prime}}=J$. Set $I_{\lambda}=K_{\lambda+1}\setminus K_{\lambda}$ and $J_{\lambda}=L_{\lambda+1}\setminus L_{\lambda}$ for every $\lambda<\lambda^{\prime}$. Then
$\bigoplus\limits_{i\in K_{\lambda+1}}P_{i}=\bigoplus\limits_{j\in L_{\lambda+1}}Q_{j}\quad\text{and}\quad\bigoplus\limits_{i\in K_{\lambda}}P_{i}=\bigoplus\limits_{j\in L_{\lambda}}Q_{j}.$
Consequently, $\bigoplus\limits_{i\in I_{\lambda}}P_{i}\cong\bigoplus\limits_{j\in J_{\lambda}}Q_{j}$ for each $\lambda\in\Lambda$. ∎
Remark 2.3.
If we set $M=R$, then $\mathrm{Add}(M)$ is the class of all projective modules. The above theorem recovers Trlifaj’s characterization of the categoricity for projective modules (see [9, Proposition 1.1]).
Recall that a class $\mathcal{D}$ of $R$-modules is called ($\kappa$-)deconstructible if there is a cardinal $\kappa$ such that every $R$-module $M$ in $\mathcal{D}$ is a transfinite extensions of modules from $\mathcal{C}$ consisting of $<\kappa$-presented modules, that is, $R$-modules $M$ that possess a $\mathcal{C}$-filtration, i.e., an increasing chain of submodules, $\mathcal{M}=(M_{\alpha}\mid\alpha\leq\sigma)$, such that $M_{0}=0$, $M_{\sigma}=M$, $M_{\alpha}=\bigcup_{\beta<\alpha}M_{\beta}$ for each limit ordinal $\alpha\leq\sigma$, and $M_{\alpha+1}/M_{\alpha}$ is isomorphic to an $R$-module in $\mathcal{C}$ for each $\alpha<\sigma$.
It was proved in [9, Theorem 2.10] that a deconstructible class $\mathcal{D}$ of $R$-modules has the categoricity property in a big cardinal is equivalent to its categoricity in a tail of cardinals (More precise expression can be found in [7, Theorem 4.1]).
Note that any deconstructible class is closed under extensions, and $\mathrm{Add}(M)$ is not closed under extensions in general. So $\mathrm{Add}(M)$ is not deconstructible in general. Hence, our main result Theorem 2.2 is not contained in [9, Theorem 2.10] or [7, Theorem 4.1].
Corollary 2.4.
Let $R$ be a ring, and $M$ be an $R$-module which contains a direct summand isomorphic to $R$ and can be generated by $\leq\aleph$ elements. Then the following statements are equivalent:
  1. (1)
    $\mathrm{Add}(M)$ is $\kappa$-categorical in all $\kappa>\|R\|+\aleph+\aleph_{0}$.
  2. (2)
    $\mathrm{Add}(M)$ is $\kappa$-categorical in some $\kappa>\|R\|+\aleph+\aleph_{0}$.
  3. (3)
    Every $R$-module of cardinal $\kappa$ in $\mathrm{Add}(M)$ is free for all $\kappa>\|R\|+\aleph+\aleph_{0}$.
  4. (4)
    Every $R$-module of cardinal $(\|R\|+\aleph+\aleph_{0})^{+}$ in $\mathrm{Add}(M)$ is free.
Proof.
If $M$ is an $R$-module which contains a direct summand isomorphic to $R$, then we can take $M$-free to be $R$-free in the proof of Theorem 2.2. ∎
Recall that an $R$-module is semisimple if it can be expressed as a direct sum of its simple submodules. The author in [5] obtained that the categoricity of semisimple modules in a big cardinal is equivalent to its categoricity in a tail of cardinals. Then he asked how can give a natural ring theoretic property for it (see [5, Question 4.1]). Now, we give an answer to the case of semisimple modules of this question.
Theorem 2.5.
Let $R$ be a ring, $S\SS$S the class of all semisimple $R$-modules and $S$ the direct sum of all simple modules non-isomorphic to each other. Then the following statements are equivalent:
  1. (1)
    $S\SS$S is $\kappa$-categorical in all $\kappa>\|R\|+\aleph_{0}$.
  2. (2)
    $S\SS$S is $\kappa$-categorical in some $\kappa>\|R\|+\aleph_{0}$.
  3. (3)
    Every semisimple module of cardinal $\kappa$ is $S$-free for all $\kappa>\|R\|+\aleph_{0}$.
  4. (4)
    Every semisimple module of cardinal $(\|R\|+\aleph_{0})^{+}$ is $S$-free.
  5. (5)
    $R$ admits a unique simple module up to isomorphism.
Proof.
If we set $M$ to be the direct sum of all simple $R$-modules non-isomorphic to each other, then $S\mathrm{Add}(M)=\SS$S. So the equivalence of (1)(4)(1)-(4)( 1 ) - ( 4 ) follows by Theorem 2.2.
(5)(4)(5$\Rightarrow$(4)( 5 ) ⇒ ( 4 ) Suppose $R$ admits a unique simple module up to isomorphism. Then every semisimple $R$-module is $S$-free.
(1)(5)(1$\Rightarrow$(5)( 1 ) ⇒ ( 5 ) On contrary, there are two non-isomorphic simple $R$-modules $S_{1}$ and $S_{2}$. Then $S_{1}^{(\kappa)}\cong S_{2}^{(\kappa)}$ for all $\kappa>\|R\|+\aleph_{0}$, which is impossible. ∎
Recall that an $R$-module $P$ is said to be a pure-projective module if any pure exact sequence $0\rightarrow A\rightarrow B\rightarrow P\rightarrow 0$ ending at $P$ splits, or equivalently, if given any pure epimorphism $g:M\rightarrow N$ and any $R$-homomorphism $f:P\rightarrow N$, there is an $R$-homomorphism $h:P\rightarrow M$ such that the following diagram is commutative
[…]
A ring $R$ is called pure-semisimple if every $R$-module is pure-projective.
It is well known that an $R$-module is pure-projective if and only if it is a direct summand of a direct sum of finitely presented $R$-modules. So one can easily deduce the following well-known result by Lemma 2.1.
Corollary 2.6.
Any pure-projective module is a direct sum of countably generated pure-projective modules.
Theorem 2.7.
Let $R$ be a ring and $P\mathcal{P}roj$ the class of all pure-projective $R$-modules. Then the following statements are equivalent:
  1. (1)
    $P\mathcal{P}roj$ is $\kappa$-categorical in all $\kappa>\|R\|+\aleph_{0}$.
  2. (2)
    $P\mathcal{P}roj$ is $\kappa$-categorical in some $\kappa>\|R\|+\aleph_{0}$.
  3. (3)
    Every pure-projective module of cardinal $\kappa$ is free for all $\kappa>\|R\|+\aleph_{0}$.
  4. (4)
    Every pure-projective module of cardinal $(\|R\|+\aleph_{0})^{+}$ is free.
  5. (5)
    $R\cong M_{n}(D)$ where $D$ is a division ring and $n\geq 1$.
Proof.
If we set $M$ to be the direct sum of all finitely presented $R$-modules non-isomorphic to each other, then $\mathrm{Add}(M)=P\mathcal{P}roj$. So the equivalence of (1)(4)(1)-(4)( 1 ) - ( 4 ) follows by Corollary 2.4.
(5)(1)(5$\Rightarrow$(1)( 5 ) ⇒ ( 1 ) Suppose $R\cong M_{n}(D)$ where $D$ is a division ring and $n\geq 1$. Then every $R$-module is free, and so the result follows by Corollary 2.4.
(1)(5)(1$\Rightarrow$(5)( 1 ) ⇒ ( 5 ) Suppose $P\mathcal{P}roj$ is $\kappa$-categorical in all $\kappa>\|R\|+\aleph_{0}$. Let $P$ be any nonzero pure-projective $R$-module. Then $P^{((\|R\|+\aleph_{0})^{+})}$ is pure-projective with $\|P^{((\|R\|+\aleph_{0})^{+})}\|\geq(\|R\|+\aleph_{0})^{+}$ and thus free by Corollary 2.4. Hence $P$ is a projective $R$-module. So every $R$-module is flat, that is, $R$ is a von Neumann regular ring. We claim that every idempotent in $R$ is trivial. Indeed, let $\{e_{\lambda}\mid\lambda\in\Lambda\}$ be a family of nonzero orthogonal idempotents in $R$. Let $\kappa>\|R\|+\aleph_{0}$. Then $(e_{\lambda}R)^{(\kappa)}$ is a pure-projective $R$-module whose cardinal is $\kappa$. Hence $(e_{\lambda}R)^{(\kappa)}\cong R^{(\kappa)}$. Consequently, $e_{\lambda}=1$. It follows by [2, Corollary 2.16] that $R$ is a semisimple ring. Consequently, $R\cong M_{n_{1}}(D_{1})\times\cdots\times M_{n_{k}}(D_{k})$ where each $D_{k}$ is a division ring and each $n_{k}\geq 1$. Since different $M_{n_{i}}(D_{i})^{(\kappa)}$s as semisimple $R$-modules have the same cardinal $\kappa$ but not isomorphic to each other, so $k=1$. Hence we have $R\cong M_{n}(D)$ where $D$ is a division ring and $n\geq 1$. ∎
Conflict of interest.
The author declares that he has no conflict of interest.
Data Availability.
My manuscript has no associated data.